# Visualizationary: Automating Design Feedback for Visualization Designers using LLMs

### v1

@@ 7 Discussion
The results from this study showed three things: that (1) a “vanilla” LLM, when provided with appropriate symbolic metrics and a “cheat sheet” of visualization design knowledge, can provide meaningful and actionable feedback on flaws in a visualization and how to improve them; that (2) even experienced visualization designers can benefit from this feedback; and that (3) the quality of visualization feedback is strongly dependent on the metrics extracted from the visualization. Below we discuss the potential of utilizing the system for novice designers. Then we enumerate limitations and avenues for future work.
7.1 On Visualization Design Feedback
@@
Our main goal for Visualizationary was to help novice designers create visualizations. Our results indeed show that our tool is capable of detecting problems and providing relevant feedback according to existing design knowledge. However, it is also clear that participants with intermediate and expert knowledge used Visualizationary more effectively than novices. Based on comments from novice participants during the post-study interview, we speculate that automation and lack of fundamental design skills could be contributing factors.
@@
By way of explanation, note that the tendency of novice users to rely on automated visualization suggestions, as offered by chart-authoring tools such as Microsoft Excel, contributes to this problem. This reliance caused some novices to lose their manual chart design skills (or never to acquire them in the first place). P12, who obtained a professional certificate for Microsoft Excel nearly a decade ago, said, “the auto-generated charts are usually of good quality, so I don’t feel the need to alter them.” P11 shared similar comments. While automated tools have simplified visualization creation, their influence may have affected our results in Table III.
Novices also have a harder time acting upon suggestions because they lack fundamental knowledge for visualization design. Many novice participants mentioned that when provided explanations and their corresponding suggestions, they still perceived it as high-level feedback; they preferred more actionable feedback where the impact of the suggested changes were clearly visible. In contrast, most participants with intermediate and expert levels understood the feedback well and used it to improve their work. P3, from the intermediate group, said, “it was a good chance to identify that what I learned in textbooks really helps.” In fact, two experts (P1 and P9) noted that the lack of specific and actionable feedback could pose challenges for novices.
On the other hand, the ineffable purpose of design is to bring novel creation forth. While Visualizationary may currently fall short in providing detailed solutions, there is also danger in hampering the designer’s creativity if suggestions become too prescriptive. The balance may need to shift depending on the experience of the designer, especially as the designer grows in experience. We leave this issue as future work.
7.2 Limitations and Future Work
While our work on Visualizationary in this paper builds on an entire stack of prior work, including the Scanner Deeply [60] virtual eye tracker and the Perceptual Pat [61] design tool, there are still weaknesses and potential shortcomings in our tool.
@@ 7.2 Limitations and Future Work
Hallucination. Despite their capabilities, LLMs may produce nonsensical or unfaithful content, known as hallucinations [33]. Such hallucinations can lead to feedback that contradicts the designer’s original goals or does not improve the visualizations. This behavior potentially harms Visualizationary because users build trust in the models powering our tool when most of their outputs are truthful. Recent work proposes defensive methods to address hallucinations, e.g., incorporating a user’s feedback into the model’s fine-tuning [52, 35] and characterizing self-contradiction in model-generated texts [48, 43]. We leave the incorporation of these methods for future work.
Privacy risks. If we use commercial services built on LLMs, there is a data privacy concern about the user data [27]. A company may collect data—such as visualizations (images) and queries (text)—that a user provides to our tool and uses them to fine-tune their models further to improve the quality of their services. Later on, a motivated adversary can perform a series of privacy attacks on the fine-tuned models, such as membership/property inference attacks [25, 10] or sensitive data-extraction attacks [11, 12]. However, this only happens if our tool is based on LLMs hosted by a third-party. We can mitigate this concern by using only local LLMs, e.g., by utilizing open-source models, such as Vicuna [14] and LLaMA [65].
Future Work. Visualizationary paves the way for several new directions for future work on automating feedback for visualization design. First, a direct approach is to add filters to analyze visualization design for improved representation of a virtual human visual system. “Will chartjunk in visualizations be identified as an object, or as a data representation by the viewer?” Answering such questions would help designers by providing more detailed, thorough feedback about their visualization designs. Another avenue for research might involve investigating the best methods—such as reports, conversational UIs that clarify vague suggestions textually, or tools offering visual examples—to deliver visualization feedback to various user groups.
An even more ambitious approach would be to expand the role of augmenting visualization design through feedback not just confined to iterative design process, but to every aspect of visualization design, including early design. We believe that the rapid development of text-to-image large language models could offer feedback that enhances designers’ creativity and expressiveness based on their intentions.
@@ 8 Conclusion
We have presented Visualizationary, a tool that leverages LLMs and visualization guidelines to provide design feedback for visualization designers. Visualizationary supports an automated visualization design workflow called ACGT (analyze-clarify-guide-track). We created a web-based interface to evaluate the effectiveness of Visualizationary and engaged 13 visualization designers of different seniority in a longitudinal design task. During this study, we conducted pre- and post-study interviews to further understand the designer’s experience using the tool. Finally, we asked three expert evaluators to assess their resulting designs. Overall, our results show that an off-the-shelf LLM can indeed provide high-quality and actionable feedback for novice, intermediate, and expert users alike on how to design a visualization.
@@ Acknowledgments
This work was partly supported by grant IIS-1901485 from the U.S. National Science Foundation (Shin), a Google Faculty Research Award (Hong), and Villum Investigator grant VL-54492 by Villum Fonden (Elmqvist). Any opinions, findings, and conclusions or recommendations expressed here are those of the authors and do not necessarily reflect the views of the funding agencies.

### v3

@@ 7 Discussion
The results from this study showed three things: that (1) a “vanilla” LLM, when provided with appropriate symbolic metrics and a “cheat sheet” of visualization design knowledge, can provide meaningful and actionable feedback on flaws in a visualization and how to improve them; that (2) even experienced visualization designers can benefit from this feedback; and that (3) our ACGT framework is effective for providing feedback for communicative visualizations. Below we discuss issues on visualization design feedback, as well as the limitations of our work.
7.1 Visualization Design Feedback
Bridging design contexts. Although we designed our feedback primarily to enhance clarity in visualizations, it is equally important to consider each designer’s unique goals and intentions. Some participants diverged from the system’s advice because it did not align with their specialized needs (see § 6.3). This underscores that relying solely on general “best practices” can overlook the nuances of personal or domain-specific objectives.
Balancing clear, communicative design with each user’s specialized goals remains a significant challenge. On one hand, emphasizing universal guidelines can neglect individual contexts; on the other, tailoring feedback for every nuanced purpose requires more extensive modeling of the designer’s intent. We see this as an important avenue for future work, where feedback systems can be refined to address both communicative effectiveness and the broader range of a designer’s motivations.
Human-like visualization feedback. AI models can help designers assess whether a visualization is “good enough” by applying best practices and accessibility guidelines. Systems like Visualizationary use automated checklists to approximate human perception, for instance by evaluating color vision deficiency and visual saliency. One way to enhance this guidance is to add more filters tailored to designers’ needs. To further validate a design or create domain-specific conditions, designers can set benchmarks (e.g., speed of comprehension) to measure success.
Nevertheless, visualization design is context-dependent and subjective, as it involves aesthetic considerations and domain-specific requirements. Can we develop a subjective design evaluator that declares a design “complete” and provides feedback in a style akin to a human visualization designer? This would not only guide users to general feedback, but also provide design feedback that would be tailored to their owners’ styles. Achieving this would require extensive knowledge of best practices, accessibility standards, and real user feedback. As a possible future direction, we plan to explore how effectively LLM-driven agents could contribute to building such models.
@@ 7.1 Visualization Design Feedback
Pitfalls of automated chart authoring. One important goal for Visualizationary was to help novice designers create visualizations. Our results indeed show that our system is capable of detecting problems and providing relevant feedback according to existing design knowledge. However, it is also clear that participants with intermediate and expert knowledge used Visualizationary more effectively than novices. Based on comments from novice participants, we speculate that automation and lack of fundamental design skills could be contributing factors.
@@
By way of explanation, note that the tendency of novice users to rely on automated visualization suggestions, as offered by chart-authoring tools such as Microsoft Excel, contributes to this problem. This reliance caused some novices to lose their manual chart design skills (or never to acquire them in the first place). N2, who obtained a professional certificate for Microsoft Excel nearly a decade ago, said, “the auto-generated charts are usually of good quality, so I don’t feel the need to alter them.” N1 shared similar comments. While automated tools have simplified visualization creation, their influence may have affected our results in Table IV.
7.2 Limitations
Avenues for future improvements. Our system is designed to enable a variety of future research avenues. To begin with, it can accommodate additional filters—for example, detecting charts that do not begin at zero (as in N2’s case) and introducing linting-like features to ensure consistency. Second, accurately predicting a saliency map for an overview task is challenging, as saliency varies by task and does not always align with overview-oriented goals. This presents the challenge of predicting saliency based on the user’s intent [68]. Third, more advanced interpretability for each filter could yield more actionable feedback for designers. For instance, refining saliency-based methods may help pinpoint issues more precisely, and automating which part of the visualization the feedback refers to would streamline problem identification. Fourth, enhancing our system’s ability to deliver user-friendly and clear feedback further expands its potential by helping designers better understand the feedback. For example, by going beyond the “if, then” structure in our design guidelines (as noted in § 4.5) and exploring alternative tones, we can improve how feedback is delivered. We can also refine our experimental procedures by adopting more neutral language (e.g., “What was your overall impression of the system?”) rather than “Did you like the system?” to mitigate bias. While we cannot address every challenge at once, we leave these avenues open for continued exploration.
@@ 7.2 Limitations
Hallucination. Despite their capabilities, LLMs may produce nonsensical or unfaithful content, commonly referred to as hallucinations [31]. Hallucinations may lead to feedback that contradicts the designer’s original goals or fails to improve the visualizations. This issue poses a potential risk to Visualizationary, as users build trust in the models powering our system when most of their outputs are consistently reliable. Recent work has proposed defensive methods to mitigate hallucinations, such as incorporating a user’s feedback into the model’s fine-tuning [33] and detecting self-contradiction in model-generated texts [46]. Incorporating these methods remains an avenue for future work.
Privacy risks. Because our system are built on LLMs, one can be concerned about the associated privacy risks, such as membership inference or data extraction attacks [11]. User data may be intentionally (or unintentionally) leaked to an adversary with access to these models [27]. The concerns primarily stem from two aspects: (1) user data being shared with commercial chat-based services and (2) the fine-tuning of these models on user data. However, the first concern can potentially be mitigated by avoiding the use of third-party hosted LLMs. We can entirely prevent our system from accessing external services by employing local open-source models. Second, we do not fine-tune the models we use on user data, ensuring that user data cannot be memorized by our LLMs or leaked by malicious actors with access to the models.
@@ 8 Conclusion
We have presented Visualizationary, a system that leverages LLMs and visualization guidelines to provide design feedback for visualization designers. Visualizationary supports an automated visualization design workflow called ACGT (analyze-clarify-guide-track). We created a web-based interface to evaluate the effectiveness of Visualizationary and engaged 13 visualization designers of different seniority in a longitudinal design task. During this study, we conducted pre- and post-study interviews to further understand the designer’s experience using the system. Finally, we asked three expert evaluators to assess their resulting designs. Overall, our results show that an off-the-shelf LLM can indeed provide high-quality and actionable feedback for novice, intermediate, and expert users alike on how to design a visualization.
@@ Acknowledgments
We thank anonymous reviewers for their feedback. This work was partly supported by grant IIS-1901485 from the U.S. National Science Foundation (Shin), a Google Faculty Research Award (Hong), and Villum Investigator grant VL-54492 by Villum Fonden (Elmqvist). Any opinions, findings, and conclusions or recommendations expressed here are those of the authors and do not necessarily reflect the views of the funding agencies.